### v2

@@ Acknowledgements
The authors are thankful to Lorenzo Federico for interesting discussions. SG and FS were partially supported by the project “CODE – Coupling Opinion Dynamics with Epidemics”, funded under PNRR Mission 4 “Education and Research” - Component C2 - Investment 1.1 - Next Generation EU “Fund for National Research Program and Projects of Significant National Interest” PRIN 2022 PNRR, grant code P2022AKRZ9.
Appendix A Maximum entropy null models for bipartite networks
A.1 Bipartite Configuration Model
The framework described in Section 2.1 can be adapted to bipartite networks. In particular, in the following we will use the degree sequences of nodes of both layers as constraints: in this case, the null-model takes the name of Bipartite Configuration Model (or BiCM [40]).
@@ A.1 Bipartite Configuration Model
Let us call the two layers of a bipartite network $\top$ and $\bot$ and let $N_{\top}$ and $N_{\bot}$ be their dimensions; nodes belonging to the $\top$-⊤ -layer will be denoted by Latin indices, while the ones belonging to the $\bot$-⊥ -layer will be denoted by Greek ones. Any bipartite network is completely described by its biadjacency matrix $\mathbf{B}$, i.e. a $N_{\top}\times N_{\bot}$ matrix whose entries $b_{i\alpha}=1$ if nodes $i\in\top$ is connected to $\alpha\in\bot$ and $b_{i\alpha}=0$ otherwise. The degree of a node $i\in\top$ ($\alpha\in\bot$), i.e. the number of connections that the nodes have, is simply $k_{i}=\sum_{\alpha}b_{i\alpha}$ ($h_{\alpha}=\sum_{i}b_{i\alpha}$).
@@
Following the procedure sketched in Section 2.1 returns that the probability of observing the generic (bipartite) graph $G_{\text{Bi}}$ is
$\begin{split}P(G_{\text{Bi}})=&\prod_{i,\alpha}\dfrac{e^{-(\theta_{i}+\eta_{\alpha})b_{i\alpha}(G_{\text{Bi}})}}{1+e^{-(\theta_{i}+\eta_{\alpha})}}\end{split}$ (3)
@@
where $\vec{\theta}$ and $\vec{\eta}$ are the Lagrangian multipliers associated with the degree sequences on, respectively, layers $\top$ and $\bot$. Eq. (3) indicates that probabilities per graph factorize in terms of independent probabilities per link $p_{i\alpha}=\frac{e^{-(\theta_{i}+\eta_{\alpha})}}{1+e^{-(\theta_{i}+\eta_{\alpha})}}$. To get the numerical values of the Lagrangian multipliers, we can impose the likelihood maximization [38].
In the present analysis, we used the python module bicm to calculate the numerical values of the probabilities: bicm, described in Ref. [49] as a submodule of NEMtropy, is also a standalone module.
A.1.1 BiCM validation of the monopartite projection of a bipartite network
@@
Using BiCM as a benchmark permits to validate the co-occurrences in the real network [50]. In fact, observed co-occurrences are simply the “projection” of the information contained in the bipartite network into one of the two layers, say $\top$:
@@
(in Ref. [40] co-occurrences were also called V-motifs since, if the two layers of the bipartite networks are drawn horizontally, they sketch a “V” between the layers). Since in BiCM the probabilities per link are independent, then
@@
In practice, the entire BiCM distribution of each $V^{ij}$ can be calculated: they are Poisson-Binomial distributions, i.e. the extension of Binomial distribution in which each event has a different probability. Calculating exactly the Poisson-Binomial distribution is computationally costly. It was shown, that when probabilities are relatively small –as in our case–, the Poisson-Binomial distribution can be finely approximated with a Poisson distribution [51]. Therefore, it is possible to efficently calculate the p-value associated with each observed co-occurrence $(V^{ij})^{*}$. Since the number of different co-occurrences is $\binom{N_{\top}}{2}$, we need multiple hypothesis testing. In the present manuscript, we implemented the False Discovery Rate (FDR,  [46]) since it controls the rate of False Positives. Again, also the validated projection used in the present manuscript was calculated using the python module bicm.
A.2 Bipartite Directed Configuration Model
Let us now consider the case in which we have a bipartite directed network and we want our model to include the information about the in- and out-degrees of nodes on both layers. It can be shown that the probability per bipartite directed graph $G_{\text{BiD}}$ factorises in terms of the BiCM probability per two undirected bipartite graphs $G_{\text{Bi}}^{\top\rightarrow\bot}$ and $G_{\text{Bi}}^{\top\leftarrow\bot}$, describing, respectively, all edges from layer $\top$ to layer $\bot$, and vice versa [41, 18]:
@@ A.2 Bipartite Directed Configuration Model
The numerical values of the $\theta_{i}^{\text{out}}$, $\theta_{i}^{\text{in}}$, $\eta_{\alpha}^{\text{out}}$ and $\eta_{\alpha}^{\text{in}}$ can be obtained through the maximization of the likelihood [38].
A.2.1 BiDCM validation of the monopartite projection of a bipartite network
Following the procedure described in Appendix A.1.1, we can provide a statistical validation of the monopartite multilink directed network obtained by projecting, say, on layer $\top$, a bipartite directed network. The number of edges from node $i$ to node $j$ in the projection is
@@ A.2.1 BiDCM validation of the monopartite projection of a bipartite network
A.2.2 The application of BiDCM-validated projection to the bipartite network of users and posts
@@
in this case, the Bipartite Configuration Model reduces to a Bipartite Partial Configuration Model [50]. Imposing that the constraints, i.e. the number of tweets authored by each user and the fact that each post has a single author, returns
@@
where we used that the degree sequence is constrained in $G_{\text{Bi}}^{\top\leftarrow\bot}$. The result in Eq. (6) is remarkable: the expected value $\langle\mathcal{V}^{ij}\rangle$ can be easily calculated without knowing explicitly the value of the Lagrangian multipliers. Therefore, p-values against BiDCM of the observed $\mathcal{V}^{ij}$ can be easily calculated, and further statistically validated using the FDR procedure described above [46].
A.2.3 Limits of applicability
Besides the simplicity of Eq. (6), there is an important issue, already mentioned for the present application in Ref. [18]. In the null model implemented above, constraints are satisfied on average. Therefore, there are configurations in the ensemble that are not physical, i.e. posts that may have either more than a single author or none. Since both non-physical configurations, i.e. more authors or none, contribute to calculating the p-values in opposite directions, their final effect is hard to predict. Nevertheless, in light of the present results and the ones of previous research, the agreement with the ground truth seems to indicate that, at least in the present application, the final effect of such an approximation is limited. Such an issue is going to be the target of further research.
Appendix B The V-measure for clustering similarity
Let us define, respectively the homogeneity and the completeness of $Q$ with respect to $P$ as
$h(P,Q)$ $=1-\frac{S(P\mid Q)}{S(P)}$
$c(P,Q)$ $=1-\frac{S(Q\mid P)}{S(Q)}.$
Note that $h=1$ if and only if the clusters obtained with $Q$ are sub-clusters of those obtained with $P$, whereas $c=1$ if and only if the clusters obtained with $Q$ are super-clusters of those obtained with $P$. It can be shown that the V-measure $\mathrm{VM}_{\beta}(P,Q)$ is the harmonic mean of $h$ and $c$, weighted by the parameter $\beta$ [31]:
$\mathrm{VM}_{\beta}(P,Q)=\frac{(1+\beta)h(P,Q)c(P,Q)}{\beta h(P,Q)+c(P,Q)}.$
It is worth noting that the VM of two partitions does not decrease linearly with the number of elements on which the two partitions disagree: the $\mathrm{VM}_{1}$ between two partitions obtained with the same algorithm is generally not greater than 0.9; if we consider two partitions that differ by only a small percentage of the network that is randomly relabeled, their $\mathrm{VM}_{1}$ is about 0.8 with $\sim$5of the network relabeled, and decreases to about 0.6 when $\sim$10-13of the network has been relabeled (cfr. Fig. 9 in Appendix B). Based on our tests, we deduce that a reasonably good $\mathrm{VM}_{1}$ between two partitions obtained with different methods/algorithms lies in the range $[0.6-0.9]$.
Fig. 9 shows the average $\mathrm{VM}_{1}$ between some partition of the Elections dataset and a noisy version of the same partition obtained randomly relabeling part of the users. We consider two choices for the initial labeling: 5 equally sized random communities and the communities obtained running the Louvain algorithm. In both cases, the $\mathrm{VM}_{1}$ rapidly decreases to about 0.8 when $\sim$5of the network has been relabeled, and to 0.6 when $\sim$10-13of the network has been relabeled. We also show that two different runs of the Louvain algorithm generate partitions whose $\mathrm{VM}_{1}$ is about 0.9, on average.
Appendix C Additional analyses
In the following, we report additional figures and tables that have been left out of the main sections. For the sake of readability and clarity, all discussions can be found in the respective captions.
Table 3: Mean and standard deviation of the execution time, expressed in seconds, of the different steps of $\mathrm{BiDC}_{V}$ on the Elections dataset, the largest dataset considered in this paper. These timings highlight that the cost of $\mathrm{BiDC}_{V}$ is dominated by the label propagation algorithm presented in Ref. [42].

### v3

@@ Discussion
Author’s contributions
S.G. and F.S. conceived and designed the study. F.S. and A.M. collected the data. S.G., F.S. and A.M. performed the modeling and data analysis with input from G.C. S.G and A.M. contributed to data preprocessing and visualization. S.G. and F.S. interpreted the results and drafted the manuscript. All authors contributed to revisions and approved the final version of the manuscript.
@@ Acknowledgements
The authors are thankful to Lorenzo Federico for interesting discussions.
Funding
@@
S.G. and F.S. were partially supported by the project “CODE – Coupling Opinion Dynamics with Epidemics”, funded under PNRR Mission 4 “Education and Research” - Component C2 - Investment 1.1 - Next Generation EU “Fund for National Research Program and Projects of Significant National Interest” PRIN 2022 PNRR, grant code P2022AKRZ9.
Appendix A Maximum Entropy Null Models
We start from a real network $G^{*}$ (in the following, all quantities related to the observed network will be denoted by an asterisk *). Firstly, let us define a set $\mathcal{G}$ – the ensemble– including all possible graphs having the same number of nodes $N$ as in $G^{*}$ [^1] Since the number of nodes is fixed in the randomization, we drop the asterisk * from $N$.. Let us identify a vector of quantities $\vec{C}$ that are fundamental for the description of the real system. A proper benchmark for $G^{*}$ would be completely random, but for the information contained in $\vec{C}(G^{*})$: in this way it is possible to highlight all properties of $G^{*}$ that are not trivially due to $\vec{C}$. Translating this rationale in maths, we need to perform a constrained maximization of the Shannon entropy (to have a maximally random benchmark) associated to the ensemble, where the constraints are exactly $\vec{C}(G^{*})$. In this sense, the (ensemble) average $\langle\vec{C}\rangle$ will be set exactly to $\vec{C}(G^{*})$.
If $S$ is the Shannon entropy, defined as
$S=-\sum_{G\in\mathcal{G}}P(G)\ln P(G),$
$P(G)=\frac{e^{-\vec{\theta}\cdot\vec{C}(G)}}{Z(\vec{\theta})},$ (1)
$\langle\vec{C}\rangle=\vec{C}(G^{*}).$ (2)
It can be shown that the condition in Eq. (2) can also be obtained through the maximization of the (log-)likelihood [Garlaschelli2008].
The framework described above is particularly flexible and can be adapted to different kinds of networks. In the following, we are interested in its bipartite versions for bipartite undirected and bipartite directed networks, i.e. respectively, the Bipartite Configuration Model (or BiCM [Saracco2015]) and the Bipartite Directed Configuration Model (or BiDCM [vanLidth2019, Becatti2019d]). In both cases, the constraints imposed in the entropy maximisation are, respectively, the undirected and directed degree sequences.
A.1 Bipartite Configuration Model (BiCM)
@@
Given a bipartite network with layers $\top$ (e.g., users) and $\perp$ (e.g., tweets), let $N_{\top}$ and $N_{\bot}$ be their dimensions. Nodes belonging to the $\top$-layer will be denoted by Latin indices, while the ones belonging to the $\bot$-layer will be denoted by Greek ones. Any bipartite network is completely described by its biadjacency matrix $\mathbf{B}$, i.e. a $N_{\top}\times N_{\bot}$ matrix whose entries $b_{i\alpha}=1$ if nodes $i\in\top$ is connected to $\alpha\in\bot$ and $b_{i\alpha}=0$ otherwise. The degree of a node $i\in\top$ ($\alpha\in\bot$), i.e. the number of connections that the nodes have, is simply $k_{i}=\sum_{\alpha}b_{i\alpha}$ ($h_{\alpha}=\sum_{i}b_{i\alpha}$).
@@
The Bipartite Configuration Model preserves the degree sequences $k_{i}=\sum_{\alpha}b_{i\alpha}$ and $h_{\alpha}=\sum_{i}b_{i\alpha}$. Following the procedure sketched above, the probability of observing a graph $G_{\text{Bi}}$ in the ensemble is:
$P(G_{\text{Bi}})=\prod_{i,\alpha}\frac{e^{-(\theta_{i}+\eta_{\alpha})b_{i\alpha}(G_{\text{Bi}})}}{1+e^{-(\theta_{i}+\eta_{\alpha})}},$
@@
where $\vec{\theta}$ and $\vec{\eta}$ are the Lagrangian multipliers associated with the degree sequences on, respectively, layers $\top$ and $\bot$.
The corresponding link probabilities are:
$p_{i\alpha}=\frac{e^{-(\theta_{i}+\eta_{\alpha})}}{1+e^{-(\theta_{i}+\eta_{\alpha})}}.$
To get the numerical values of the Lagrangian multipliers, we can impose the likelihood maximization [Garlaschelli2008].
A.1.1 Validated Projection under BiCM
@@
Using BiCM as a benchmark permits to validate the co-occurrences in the real network [Saracco2017]. In fact, observed co-occurrences are simply the “projection” of the information contained in the bipartite network into one of the two layers, say $\top$:
@@
(in Ref. [Saracco2015] co-occurrences were also called V-motifs since, if the two layers of the bipartite networks are drawn horizontally, they sketch a “V” between the layers). Since in BiCM the probabilities per link are independent, then
@@
In practice, the entire BiCM distribution of each $V^{ij}$ can be calculated: they are Poisson-Binomial distributions, i.e. the extension of Binomial distribution in which each event has a different probability. Calculating exactly the Poisson-Binomial distribution is computationally costly. It was shown, that when probabilities are relatively small –as in our case–, the Poisson-Binomial distribution can be finely approximated with a Poisson distribution [Hong2013]. Therefore, it is possible to efficently calculate the p-value associated with each observed co-occurrence $(V^{ij})^{*}$. Since the number of different co-occurrences is $\binom{N_{\top}}{2}$, we need multiple hypothesis testing. In the present manuscript, we implemented the False Discovery Rate (FDR,  [Benjamini1995]) since it controls the rate of False Positives. Again, also the validated projection used in the present manuscript was calculated using the python module bicm.
A.2 Bipartite Directed Configuration Model (BiDCM)
Let us now consider the case in which we have a bipartite directed network and we want our model to include the information about the in- and out-degrees of nodes on both layers. It can be shown that the probability per bipartite directed graph $G_{\text{BiD}}$ factorises in terms of the BiCM probability per two undirected bipartite graphs $G_{\text{Bi}}^{\top\rightarrow\bot}$ and $G_{\text{Bi}}^{\top\leftarrow\bot}$, describing, respectively, all edges from layer $\top$ to layer $\bot$, and vice versa [vanLidth2019, Becatti2019d]:
@@ A.2 Bipartite Directed Configuration Model (BiDCM)
The numerical values of the $\theta_{i}^{\text{out}}$, $\theta_{i}^{\text{in}}$, $\eta_{\alpha}^{\text{out}}$ and $\eta_{\alpha}^{\text{in}}$ can be obtained through the maximization of the likelihood [Garlaschelli2008].
A.2.1 Validated Projection under BiDCM
As done for the BiCM, we can provide a statistical validation of the monopartite multilink directed network obtained by projecting, say, on layer $\top$, a bipartite directed network. The number of edges from node $i$ to node $j$ in the projection is
@@ A.2.1 Validated Projection under BiDCM
A.2.2 BiDCM-Validated Projection for Users and Posts
@@
in this case, the Bipartite Configuration Model reduces to a Bipartite Partial Configuration Model [Saracco2017]. Imposing that the constraints, i.e. the number of tweets authored by each user and the fact that each post has a single author, returns
@@
where we used that the degree sequence is constrained in $G_{\text{Bi}}^{\top\leftarrow\bot}$. The result in Eq. (5) is remarkable: the expected value $\langle\mathcal{V}^{ij}\rangle$ can be easily calculated without knowing explicitly the value of the Lagrangian multipliers. Therefore, p-values against BiDCM of the observed $\mathcal{V}^{ij}$ can be easily calculated, and further statistically validated using the FDR procedure described above [Benjamini1995].
A.2.3 Limitations
Besides the simplicity of Eq. (5), there is an important issue, already mentioned for the present application in [Becatti2019d]. In the null model implemented above, constraints are satisfied on average. Therefore, there are configurations in the ensemble that are not physical, i.e. posts that may have either more than a single author or none. Since both non-physical configurations, i.e. more authors or none, contribute to calculating the p-values in opposite directions, their final effect is hard to predict. Nevertheless, in light of the present results and the ones of previous research, the agreement with the ground truth seems to indicate that, at least in the present application, the final effect of such an approximation is limited. Such an issue is going to be the target of further research.
A.3 Implementation
In the paper, we used the implementation of the BiCM and BiDCM available in the python module NEMtropy, described in [Vallarano2021]. bicm is also available as a standalone module. As a reference, the running time for BiDCM on the largest dataset are reported in Table 3.
Table 3: Mean and standard deviation of the execution time, expressed in seconds, of the different steps of $\mathrm{BiDC}_{V}$ on the Elections dataset, the largest dataset considered in this paper. These timings highlight that the cost of $\mathrm{BiDC}_{V}$ is dominated by the label propagation algorithm presented in [Raghavan2007a].
@@ A.3 Implementation
Appendix B V-Measure for Clustering Similarity
To assess the similarity between two ways of partitioning the same set of users, we use of the V-measure ($\mathrm{VM}_{\beta}$), a parametric measure of cluster similarity proposed in [rosenberg2007v].
$\mathrm{VM}_{\beta}(P,Q)=\frac{(1+\beta)I(P,Q)}{S(P)+\beta S(Q)}$
where $I(P,Q)$ is the mutual information between $P$ and $Q$. $VM_{\beta}$ can be interpreted as the harmonic mean, weighted by $\beta$, of homogeneity $h$ and completeness $c$:
$h(P,Q)=1-\frac{S(P|Q)}{S(P)},\quad c(P,Q)=1-\frac{S(Q|P)}{S(Q)},$
$VM_{\beta}(P,Q)=\frac{(1+\beta)hc}{\beta h+c}.$
Note that $h=1$ if and only if the clusters obtained with $Q$ are sub-clusters of those obtained with $P$, whereas $c=1$ if and only if the clusters obtained with $Q$ are super-clusters of those obtained with $P$ [rosenberg2007v].
For $\beta=1$ we have
i.e., the $\mathrm{VM}_{1}$ accounts for the mutual information between the two partitions in a way that does not depend on their entropy, being equivalent to the Variation of Information [Meila2007], normalized by the total entropy of the partitions. With respect to other information-theoretic metrics, the $\mathrm{VM}_{1}$ and the VI have the advantage of being two distances – in particular, of satisfying the triangle inequality. Further, the $\mathrm{VM}_{1}$ always takes values in $[0,1]$, making it possible to directly compare the scores obtained for different combinations of algorithms and/or datasets. By letting $\beta$ vary in $0,+\infty)$, we consider both scenarios in which one partition consists of sub-communities (e.g., political movements/trends within parties) or super-communities (e.g., political coalitions) of the other one.
It is worth noting that the VM of two partitions does not decrease linearly with the number of elements on which the two partitions disagree. Fig. 9 shows the average $\mathrm{VM}_{1}$ between some partition of the Elections dataset and a noisy version of the same partition obtained randomly relabeling part of the users. We consider two choices for the initial labeling: 5 equally sized random communities and the communities obtained running the Louvain algorithm. In both cases, the $\mathrm{VM}_{1}$ rapidly decreases to about 0.8 when $\sim$5% of the network has been relabeled, and to 0.6 when $\sim$10-13% of the network has been relabeled. We also show that two different runs of the Louvain algorithm generate partitions whose $\mathrm{VM}_{1}$ is about 0.9, on average. Based on these tests, we deduce that a reasonably good $\mathrm{VM}_{1}$ between two partitions obtained with different methods/algorithms lies in the range [0.60.9][0.6-0.9].
Appendix C Data collection and political context
Data were downloaded using Twitter/X official APIs, using a Researcher account. Data were collected using the Python module twarc2 (https://twarc-project.readthedocs.io/en/latest/twarc2_en_us/), using the search --archive command to access the archive of data.
Table 4: Description of the queries used for the data collection
Dataset Keywords Start time End time
President quirinale 2022-01-01 2022-01-31
presidente della repubblica
mattarella
Crisis crisidigoverno 2022-07-01 2022-08-01
crisigoverno
draghi
governo
Elections elezionipolitiche2022 2022-08-25 2022-10-02
elezionipolitiche22
elezioni
PD
Letta
Calenda
Azione
Renzi
ItaliaViva
FdI
Meloni
Lega
Salvini
Conte
M5S
ForzaItalia
Berlusconi
In the President dataset, “Quirinale” is the name of the Roman hill where the headquarters of the Italian President of the Republic is. “Presidente della Repubblica” is the Italian name for President of the Republic, while “Mattarella” is the surname of the president that was confirmed in early 2022.
In 2021, President Mattarella consistently declined to make himself available for a second term, despite sustained pressure from a broad range of political actors. Nonetheless, in the absence of an officially endorsed candidacy, he continued to attract a significant number of votes in successive ballots, obtaining 336 votes in the sixth round and 387 in the seventh. By 29 January, the failure of all alternative candidacies and the persistent coordination breakdown between the two major coalitions rendered the re-election of the incumbent a focal point of convergence. Following this deadlock, Mattarella accepted a renewed mandate and was re-elected with 759 votes.
In Italian political discourse, “crisi di governo” is an idiomatic expression referring to a cabinet crisis, namely the fall of the executive without a systemic constitutional breakdown. Mario Draghi was the Prime Minister at the time of the crisis.
On July 13, 2022, during a press conference, Conte, leader of the M5S, declared that his party would abstain from voting on the Aid Decree, which the government had designated as a matter of confidence. The following day, the Senate approved the decree, while M5S representatives withdrew from the chamber during the vote. Although this action did not constitute a formal withdrawal of governmental support, it was broadly interpreted as a substantial challenge to the administration’s agenda and effectively precipitated a political crisis within Draghi’s Cabinet. After strong debates in the two chambers of the Italian parliament, on July 21, Draghi definitively submitted his resignation to President Mattarella, who was compelled to accept, as the government’s loss of parliamentary support had by then become unequivocal and irreversible.
The keywords used to gather tweets for the Elections dataset include the surname of the main political party leaders (i.e. Letta, Calenda, Renzi, Meloni, Salvini, Conte, Berlusconi), as well as the name of their party (PD, Azione, Italia Viva, Fratelli d’Italia, Lega, Movimento 5 Stelle, Forza Italia); other keywords relate to the event itself.
Following the Draghi government’s collapse and resulting parliamentary deadlock, President Sergio Mattarella dissolved Parliament on July 21, 2022, calling early elections for September 25. The centre-right coalition led by Giorgia Meloni’s national-conservative Fratelli d’Italia (FdI) won an absolute parliamentary majority. Meloni was appointed Prime Minister on October 22, becoming Italy’s first female head of government.
Appendix D Experimental Details and Extended Comparisons
D.1 Community Detection Algorithms and Implementations
Community detection algorithms used:
  • Louvain, Label Propagation, Infomap from the igraph Python package.
  • Stochastic Block Model (SBM) inference via graph-tool’s minimize_blockmodel_dl().
The minimize_blockmodel_dl() function fits a stochastic block-model to the given graph by minimizing the model’s description length. To guide the inference process, we set it explicitly to search for a planted-partition, that is, exclusively for assortative communities. In addition, we refined the algorithm as recommended in graph-tool’s documentation, that is, by running 100 times the multiflip_mcmc_sweep() method with zero temperature and 100 iterations.
We ran each algorithm 100 times. Each single run of the Louvain algorithm is already obtained taking the partition with the greatest modularity out of 100 independent executions with shuffled vertices. Similarly, each run of the Label propagation algorithm is already obtained by taking, for each node, the most frequent label out of 100 independent executions.
D.2 Definition of $H$ Users
We repeated MonoDC and BiDC using:
  • Influential users ($I$): users with follower/followee ratio >1>1 and more mentions received than made.
  • Top H-index users ($H$): users with $H\geq 3$ where $H$ is the largest number such that the user has at least $H$ tweets each with at least $H$ retweets.
The threshold $H>3$ has been selected based on the elbow-method applied to the frequency of users for whom $H>h$ for different values of $h$ (cfr. Fig. 10).
D.3 Effect of Significance Thresholds
When running the $\mathrm{BiDC}_{V}$ method, we observed that only a small percentage of all edges resulted as being statistically significant. We therefore explored BiDCM validation at multiple significance levels ($\alpha=0.001$, $0.01$, $0.05$). Results, shown in Fig. 11, were stable across values, confirming robustness. Non-validated projections performed poorly.